OpenStudy (anonymous):

Find all solutions of the equation in the complex number system: x^2+4x+4=-9

jhonyy9 (jhonyy9):

\[x ^{2}+4x +13=0\]

jhonyy9 (jhonyy9):

x_1,_2= (-4+/- sqrt(16-52))/2 =(-4+/- sqrt(-36))/2 = =(-4+/-6i)/2=-2+/-3i

OpenStudy (anonymous):

thanks

jhonyy9 (jhonyy9):

my pleasure